OpenStudy (anonymous):

log_{2}(x-10)-3=log_{2}(1-x)

OpenStudy (zarkon):

\[\log_2(x-10)-3=\log_2(1-x)\] \[\log_2(x-10)-\log_2(1-x)=3\] \[\log_2\left(\frac{x-10}{1-x}\right)=3\]

OpenStudy (zarkon):

\[\frac{x-10}{1-x}=2^3\] continue to solve for x ...

OpenStudy (zarkon):

you should get x=2

OpenStudy (zarkon):

I should state here that this is only true if you are allow to work in the complex plane. clearly is you put in x=2 into the above equation you get negative values inside the log. if you are just working in R then you really have no solution. if you are allowed to work in C the 2 is the solution